The drug Lipitor is mean to reduce cholesterol and LDL cholesterol. I n clinical trials 19 out of 863 patients taking 10 mg of Lipitor daily complained of flulike symptoms. Suppose that it is known that 1.9% of patients taking competing drugs complain of flulike symptoms. Is there evidence to conclude that more than 1.9% of Liptor users experience flulike-smptoms as a side effect at α=0.05, level of significance ?

Which of the following is true In the hypothesis testing of the above problem?

1.The conclusion is do not reject

2.The p-value < α=0.01

3.The test is two tail test

4.The test statistic is z = 1.56
Answer #1

Solution:

The null and alternative hypothesis are

H0 : p = 0.019 vs Ha : p > 0.019

n = 863

x = 19

Let  \hat p be the sample proportion.

\hat p = x/n = 19/863 = 0.0220

The test statistic z is

z =   Р-р р(1-р) п

=  (0.0220 - 0.019)/\sqrt{}[0.019*(1 - 0.019)/863]

= 0.65

The Test Statistic is z = 0.65

> sign in Ha indicates that the test is "RIGHT TAILED"(One tailed right sided)

For right tailed z test ,

p value = P(Z > z) = P(Z > 0.65) = 1 - P(Z < 0.65) = 1 - 0.7422 = 0.2578

p value = 0.2578

p value > α = 0.01

The conclusion is do not reject.

Correct option is

The conclusion is do not reject